OpenStudy (anonymous):

Find solutions to the equation: xy′+2y=x3 such that y = 1 when x = 1

OpenStudy (anonymous):

is that differential equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[xy'+2y=x^3\] divide by x \[y'+2y/x=x^2\]

OpenStudy (anonymous):

\[y=u*v\] \[y=u'*v+u*v'\] Put into equation \[u'*v+u*v'+2uv/x=x^2\] \[u'*v+u*(v'+2v/x)=x^2\]

OpenStudy (anonymous):

Let \[v'+2v/x=0\] (1) then equation \[u'v=x^2\] (2)

OpenStudy (zarkon):

\[xy'+2y=x^3\] multiply by x \[x^2y'+2yx=x^4\] \[\frac{d}{dx}(x^2y)=x^4\] integrate both sides ... solve for y

OpenStudy (zarkon):

\[x^2y=\frac{1}{5}x^5+c\] \[y=\frac{1}{5}x^3+\frac{c}{x^2}\] now solve for the initial condition

OpenStudy (anonymous):

(1)\[dv/dx=-2v/x\] \[\int\limits( dv/v)=-2\int\limits dx/x\]

OpenStudy (anonymous):

\[\ln(v)=-2\ln(x)\]\[v=x^{-2}\]

OpenStudy (anonymous):

Put \[v=x^{-2}\] into (2) \[u'v=x^2\]\[u'x^{-2}=x^2\]

OpenStudy (anonymous):

\[u'=x^4\]\[u=x^5/5+C\]

OpenStudy (anonymous):

\[y=u*v=(x^5/5+C)*x^{-2}=x^3/5+C/x^2\]

OpenStudy (anonymous):

y = 1 when x = 1 Put into the solution y=1 and x=1 \[1=1/5+C\], then C=1-1/5=4/5 So solution \[y=x^3/5+4/(5x^2)\]

OpenStudy (anonymous):

Sorry, my laptop died when I was working through that, I was following until I got to: Let v′+2v/x=0

OpenStudy (anonymous):

@ Tatiana_Stebko is correct.\[y[x]= \frac{4}{5 x^2}+\frac{x^3}{5}\]\[x y'[x]+2y[x]\text{= }x^3\]\[x \left(-\frac{8}{5 x^3}+\frac{3 x^2}{5}\right)+2 \left(\frac{4}{5 x^2}+\frac{x^3}{5}\right)=x^3 \]\[\frac{-8+3 x^5}{5 x^2}+\frac{2 \left(4+x^5\right)}{5 x^2}= x^3\]\[x^3 = x^3\]

OpenStudy (anonymous):

Why can we let v′+2v/x=0

OpenStudy (zarkon):

The easiest way to do this problem is with an integrating factor.